# Explainable Framework for Swarm Intelligence Based on Fitness Landscape Features and Machine Learning

### v2

@@ 1 Introduction
Swarm Intelligence (SI) represents a fascinating and powerful field of study that draws inspiration from the collective behaviour observed in decentralized, self-organized natural systems [7]. The complex collective intelligence observed in nature, from the synchronized movements of bird flocks to the optimized foraging of ant colonies, provides the foundational inspiration for sophisticated optimization algorithms. At its core, Swarm Intelligence (SI) is characterized by a population of simple, autonomous agents that follow basic rules and interact locally with their environment and neighbors. Without any centralized guidance, these limited interactions collectively produce sophisticated global patterns and solve complex problems. The approach is fundamentally defined by principles such as decentralization, emergence [7] , adaptation, and positive feedback. Any swarm based optimization algorithm includes three fundamental components: (i) a population that evolves iteratively, (ii) a reproduction process where current population of candidate solutions generates new set of candidate solutions in the next generation, and (iii) a curated information sharing along with an efficient learning mechanism that determines next generation of candidate solutions. These algorithms are inherently scalable, adaptable, and robust, characterized by the simplicity of their individual candidate solutions, making them highly effective for solving a wide array of complex optimization problems [8, 4, 42]. The efficacy of SI algorithms is fundamentally dependent on a critical balance: the exploration of uncharted regions of the search space to discover novel solutions, and the exploitation of known promising areas to refine them. Conversely, exploitation involves refining the search around known promising solutions, leveraging existing information to converge efficiently. Achieving an optimal balance between these two dynamics is paramount for efficiently converging to acceptable solutions while simultaneously avoiding premature convergence to suboptimal local optima. This decentralized architecture is the direct source of the method’s primary strengths: robustness and adaptability [7]. Unlike traditional top-down optimization, the overall solution emerges from the bottom-up interactions of the swarm. This emergent quality makes such systems intrinsically flexible and resilient to dynamic, high-dimensional problem spaces, as the loss of any single agent does not derail the collective process. If one candidate solution fails or a local condition changes, the overall swarm can still adapt and find solutions because its intelligence is distributed, not concentrated in a single vulnerable point. This distributed intelligence makes SI algorithms particularly well-suited for problems where the environment is uncertain, dynamic, or too complex for a single, pre-programmed control mechanism [15]. Among the myriad of SI algorithms, several have gained prominence due to their effectiveness and versatility. Particle Swarm Optimization (PSO) [23] is a prominent algorithm in this category, modeled on the social dynamics of species like birds and fish. In PSO, particles (candidate solutions) traverse a search space, continually adjusting their positions based on their personal best (pbest) known position and the global best (gbest) position discovered by the entire swarm. PSO is particularly well-suited for continuous optimization problems [29], such as the crucial task of parameter tuning [21] in machine learning models. There are several other swarm based optimization algorithms such as artificial bee colony (ABC)[22], ant colony optimization (ACO) [12] and many more but our focus in this article is particle swarm optimization (PSO) which is one of the earliest swarm based optimization algorithms. Parameters play a fundamental role in the effective performance of swarm based optimization algorithms, profoundly influencing a meta-heuristic’s efficiency in solving a given decision problem [2]. The primary function of these parameters is to carefully balance the convergence and diversity of the search process, which directly translates to managing the trade-off between exploration (searching new areas of the solution space) and exploitation (refining the search around currently identified promising solutions) [48]. Suboptimal selection of these control parameters can lead to significant issues, such as premature convergence, where the swarm settles on a suboptimal solution too quickly, and an increase in computational time. The quality of the solution, the amount of computational power needed, the degree of diversity and exploration attained, and the general leadership dynamics within the swarm are all directly impacted by parameter settings [8]. Putting our focus back on PSO, its performance is notably sensitive to its control parameters, particularly the inertia weight $(w)$ and the acceleration coefficients $(c_{1},c_{2})$ [48].Throughout the search process, dynamically balancing exploration and exploitation requires adaptive strategies for these parameters. Our comprehension of PSO parameters is further improved by the following points:
@@
    Number of Iterations: While the number of iterations is often just a stopping criterion, studies suggest it has a less substantial impact on performance compared to key parameters such as inertia weight and swarm size [6].
A complex trade-off between convergence speed and the possibility of becoming trapped in local optima is frequently the result of the PSO algorithm’s intrinsic stochasticity and non-linear dynamics, which indicate that parameter settings have a disproportionately large impact on performance [48]. This makes manual tuning a challenging endeavour. Because SI algorithms (including PSO) are stochastic and their search behaviour emerges from simple rules, even minor changes in initial parameter values can result in vastly different search trajectories and outcomes, ranging from divergence to premature convergence or successful optimization. The direct ways to determine optimal parameters PSO are very limited in literature for a given problem, which may have non-linearity and interdependency. This forces researchers to adopt trial-and-error approach, often relying on intuition and extensive experimentation rather than precise calculation [14]. This fundamental challenge of parameter sensitivity is the primary driver for the development of more sophisticated, adaptive parameter strategies, as static and empirically derived parameters are rarely optimal across diverse problem landscapes in the entire optimization process [45].
@@ 1.1 Need of Explainabilty in PSO and Current Status of Research
Despite its demonstrated efficacy and widespread adoption, PSO, akin to many advanced AI algorithms, it (PSO) frequently functions as a “black box,” [59] obscuring its internal decision-making processes. This inherent lack of clarity poses significant challenges for building trust and verifying the reliability of its outcomes, particularly when applied in sensitive or critical domains like healthcare and autonomous systems. Current literature reveals a multi-faceted approach to interpreting these complex, self-organizing systems. A prominent strategy involves post-hoc explanation methods, where surrogate models—such as decision trees or linear models—are trained on the input-output data generated by the swarm to provide a global, simplified summary of solution features [30]. Complementing this, social network analysis (SNA) is used to model the dynamic communication topology of a swarm, quantifying the influence of individual agents and tracing the flow of information that leads to convergence [24]. For a more intrinsic understanding, fitness landscape analysis (FLA) characterizes the problem’s topology to explain algorithmic performance, such as justifying convergence stalls by identifying multimodal ruggedness [38]. Furthermore, visual analytics and dimensionality reduction techniques are leveraged to create interactive dashboards and project high-dimensional swarm trajectories into 2D or 3D spaces, allowing users to visually track exploration and exploitation behavior over time [61]. Despite these advances, challenges remain in standardizing evaluation metrics for explanations and developing real-time, human-in-the-loop XAI frameworks to fully bridge the gap between powerful optimization and user trust. The ambiguous nature of PSO’s configurations and hyperparameters can directly contribute to a perceived lack of reliability in the solutions generated by the algorithm. The stochastic nature of particle movements and the emergent, non-linear interactions within the swarm contribute significantly to this opacity, making it difficult for human users to understand why a specific optimal solution was found or how the swarm navigated the complex search space to converge. This apparent contradiction between PSO’s algorithmic simplicity and its behavioural opacity, reveals a deeper truth: PSO’s algorithmic rules are simple to implement, yet its collective behavior is highly complex and inscrutable. This obscurity does not arise from intricate internal structures but from the swarm’s emergent “intelligence”—the intricate interplay of local and global information sharing across a high-dimensional space. The challenge for XAI is thus that the “black box” nature of such metaheuristics stems from the non-linear dynamics of many simple agents, which confounds a clear tracing of the causal pathway to a solution. Therefore, XAI methods for PSO need to focus on analyzing collective behavior and parameter sensitivity rather than just decomposing complex architectures. In response to this challenge, the fields of Explainable Artificial Intelligence (XAI) and Interpretable Artificial Intelligence (IAI) [13] have rapidly gained prominence. Their primary objective is to develop methodologies that render AI models inherently understandable, thereby clarifying their decision-making processes and nurturing greater confidence in their results.
@@
Historically, metaheuristic optimization algorithms like PSO were primarily evaluated based on their efficiency, convergence speed, and the quality of the solutions they found. However, recent work [17] on analysing the impact of swarm topologies provides an early work in the direction of “enhancing explainability and trustworthiness” in PSO. This work is an extension of the work carried out in [17]. This signals a significant maturation within the field of computational intelligence. It indicates that achieving optimal solutions is no longer the sole criterion for algorithm success. Instead, there is a growing recognition that for AI systems to be responsibly deployed and widely adopted, especially in critical applications, they must also be transparent and understandable [3]. This reflects a broader ethical and practical demand for trustworthy AI. This paradigm shift implies a pressing need for the development of new evaluation metrics that extend beyond traditional optimization performance (e.g., solution quality, computational cost) to quantitatively assess explainability. It also suggests that future research in PSO will increasingly integrate XAI considerations from the very initial design phases of new algorithms, rather than merely applying post-hoc explanations [50]. The follwoing subsections further throws light on some key factors which have the highest impact in the interpretability and explainability of PSO.
@@ 1.1.1 Impact of Communication Topologies in PSO
The effectiveness and search behavior of PSO are profoundly influenced by its communication topology, which defines the structural pattern through which particles share information within the swarm. Common topologies include Ring, [32, 52] Star [41, 43], and Von Neumann [56, 37]. Research investigates how these distinct topologies impact critical aspects of PSO’s operation, such as information flow, diversity maintenance, convergence speed, and the delicate balance between exploration (searching new regions) and exploitation (refining known good regions) [60]. By systematically analyzing these influences, it is possible to enhance the inherent interpretability of PSO’s decision-making processes and derive practical guidelines for selecting the most suitable topology for specific optimization tasks. The explicit detailing of how different communication topologies directly affect PSO’s convergence, diversity, and exploration-exploitation balance provides a powerful form of intrinsic interpretability. This approach is powerful because it moves beyond treating PSO as an indivisible black box. Instead, it focuses on understanding and designing its core components to be inherently transparent. This suggests that modifying the algorithm’s internal architecture and interaction rules can directly contribute to its interpretability, offering a promising avenue for developing more transparent and trustworthy metaheuristics, in future.
@@ 1.1.2 Explainable Benchmarking Frameworks for PSO
The development and application of specialized frameworks, such as the adapted IOHxplainer [55], serve as crucial explainable benchmarking tools. These tools can be used to systematically investigate and quantify how different algorithmic configurations, including communication topologies, influence PSO’s performance characteristics. These frameworks are sophisticated enough to support continuous, integer, and categorical parameters, and critically, they account for hyperparameter dependencies, allowing for a detailed analysis of PSO configurations. They often integrate post-hoc XAI methods, such as SHapley Additive exPlanations (SHAP-based techniques) [36], to determine the marginal impact of each parameter on PSO’s performance, thereby significantly aiding in the interpretability of the experimental results.
@@ 1.1.3 Deeper Understanding Benchmark Problems
Understanding the behaviour of optimization algorithms on benchmark problems is not straightforward, as problem difficulty often depends on hidden structural properties of the search space. Relying only on performance metrics such as convergence speed or solution quality provides limited insight, since these results do not explain why an algorithm succeeds or fails. This is where tools like Exploratory Landscape Analysis (ELA) [39] can play a very crucial role. ELA features allow us to quantify the geometry and structure of optimization problems in a systematic manner. By measuring properties like ruggedness, modality, separability, and variable scaling, ELA helps uncover the reasons behind an algorithm’s performance. For example, highly rugged landscapes may demand more global exploration, while separable problems can be effectively handled by coordinate-wise methods. Such insights are crucial for moving beyond trial-and-error experimentation and towards principled algorithm design and selection. In essence, ELA features have potential to bridge the gap between “black-box” benchmarking and meaningful interpretation. These insights not only guide us in diagnosing problem hardness but also aid in designing adaptive and generalizable optimization strategies.
@@ 1.3 Summary of Contributions
An metaheuristic optimization algorithm’s performance is largely determined by the particular issue it is trying to solve. To figure out why some problems seem tougher for certain algorithms, we look into how algorithms and problems interact. This approach characterizes optimization problems by extracting meta-features from candidate solutions, which are typically generated using sampling techniques such as Latin hypercube sampling  [49], random sampling, or Sobol sequences [9]. These sampled solutions form the basis for computing landscape features that capture essential problem properties. For explainable analysis in machine learning, several powerful techniques exist to derive these features, these approaches provide a comprehensive framework for understanding and comparing optimization problems in a meaningful way.
@@
This work extends our earlier work on the explainability of PSO communication topologies [17], further enhancing interpretability through the integration of ELA and IOHxplainer-based analysis. The remainder of this paper is organized as follows: Section 2 presents the ELA-based visualization framework employed for benchmarking. Section 3 introduces the proposed explainability framework for PSO, followed by Section 4, which details the experimental setup and corresponding results. The influence of hyperparameters on algorithmic performance is examined in Section 5, while Section 6 outlines the data-driven, interpretable configuration learning methodology. Finally, Section 7 provides the concluding remarks and potential directions for future research.

### v3

# Explainable Swarm: A Methodological Framework for Interpreting Swarm Intelligence

@@ 1 Introduction
Swarm Intelligence (SI) represents a fascinating and powerful field of study that draws inspiration from the collective behaviour observed in decentralized, self-organized natural systems [bonabeau1999swarm, gonzalez2021addressing]. The complex collective intelligence observed in nature, from the synchronized movements of bird flocks to the optimized foraging of ant colonies, provides the foundational inspiration for sophisticated optimization algorithms. At its core, Swarm Intelligence (SI) is characterized by a population of simple, autonomous agents that follow basic rules and interact locally with their environment and neighbors. Without any centralized guidance, these limited interactions collectively produce sophisticated global patterns and solve complex problems. The approach is fundamentally defined by principles such as decentralization, emergence [bonabeau1999swarm] , adaptation, and positive feedback. Any swarm based optimization algorithm includes three fundamental components: (i) a population that evolves iteratively, (ii) a reproduction process where current population of candidate solutions generates new set of candidate solutions in the next generation, and (iii) a curated information sharing along with an efficient learning mechanism that determines next generation of candidate solutions. These algorithms are inherently scalable, adaptable, and robust, characterized by the simplicity of their individual candidate solutions, making them highly effective for solving a wide array of complex optimization problems [bonyadi2017particle, bianchi2009survey, munoz2015algorithm]. The efficacy of SI algorithms is fundamentally dependent on a critical balance: the exploration of uncharted regions of the search space to discover novel solutions, and the exploitation of known promising areas to refine them. Conversely, exploitation involves refining the search around known promising solutions, leveraging existing information to converge efficiently. Achieving an optimal balance between these two dynamics is paramount for efficiently converging to acceptable solutions while simultaneously avoiding premature convergence to suboptimal local optima. This decentralized architecture is the direct source of the method’s primary strengths: robustness and adaptability [bonabeau1999swarm]. Unlike traditional top-down optimization, the overall solution emerges from the bottom-up interactions of the swarm. This emergent quality makes such systems intrinsically flexible and resilient to dynamic, high-dimensional problem spaces, as the loss of any single agent does not derail the collective process. If one candidate solution fails or a local condition changes, the overall swarm can still adapt and find solutions because its intelligence is distributed, not concentrated in a single vulnerable point. This distributed intelligence makes SI algorithms particularly well-suited for problems where the environment is uncertain, dynamic, or too complex for a single, pre-programmed control mechanism [engelbrecht2005fundamentals]. Among the myriad of SI algorithms, several have gained prominence due to their effectiveness and versatility. Particle Swarm Optimization (PSO) [kennedy1995new, yu2025particle] is a prominent algorithm in this category, modeled on the social dynamics of species like birds and fish. In PSO, particles (candidate solutions) traverse a search space, continually adjusting their positions based on their personal best (pbest) known position and the global best (gbest) position discovered by the entire swarm. PSO is particularly well-suited for continuous optimization problems [li2021particle], such as the crucial task of parameter tuning [hossain2021machine] in machine learning models. There are several other swarm based optimization algorithms such as artificial bee colony (ABC)[karaboga2007powerful, coleto2020artificial], ant colony optimization (ACO) [dorigo2018introduction, liu2023improved] and many more but our focus in this article is particle swarm optimization (PSO) which is one of the earliest swarm based optimization algorithms. Parameters play a fundamental role in the effective performance of swarm based optimization algorithms, profoundly influencing a meta-heuristic’s efficiency in solving a given decision problem [banks2007review]. The primary function of these parameters is to carefully balance the convergence and diversity of the search process, which directly translates to managing the trade-off between exploration (searching new areas of the solution space) and exploitation (refining the search around currently identified promising solutions) [shi1998modified]. Suboptimal selection of these control parameters can lead to significant issues, such as premature convergence, where the swarm settles on a suboptimal solution too quickly, and an increase in computational time. The quality of the solution, the amount of computational power needed, the degree of diversity and exploration attained, and the general leadership dynamics within the swarm are all directly impacted by parameter settings [bonyadi2017particle]. Putting our focus back on PSO, its performance is notably sensitive to its control parameters, particularly the inertia weight $(w)$ and the acceleration coefficients $(c_{1},c_{2})$ [shi1998modified].Throughout the search process, dynamically balancing exploration and exploitation requires adaptive strategies for these parameters. Our comprehension of PSO parameters is further improved by the following points:
@@
    Number of Iterations: While the number of iterations is often just a stopping criterion, studies suggest it has a less substantial impact on performance compared to key parameters such as inertia weight and swarm size [blackwell2007particle].
A complex trade-off between convergence speed and the possibility of becoming trapped in local optima is frequently the result of the PSO algorithm’s intrinsic stochasticity and non-linear dynamics, which indicate that parameter settings have a disproportionately large impact on performance [shi1998modified]. This makes manual tuning a challenging endeavour. Because SI algorithms (including PSO) are stochastic and their search behaviour emerges from simple rules, even minor changes in initial parameter values can result in vastly different search trajectories and outcomes, ranging from divergence to premature convergence or successful optimization. The direct ways to determine optimal parameters PSO are very limited in literature for a given problem, which may have non-linearity and interdependency. This forces researchers to adopt trial-and-error approach, often relying on intuition and extensive experimentation rather than precise calculation [eiben2011parameter]. This fundamental challenge of parameter sensitivity is the primary driver for the development of more sophisticated, adaptive parameter strategies, as static and empirically derived parameters are rarely optimal across diverse problem landscapes in the entire optimization process [pedersen2010tuning].
@@ 1.1 Need of Explainabilty in PSO and Current Status of Research
Despite its demonstrated efficacy and widespread adoption, PSO, akin to many advanced AI algorithms, it (PSO) frequently functions as a “black box,” [yang2025meta] obscuring its internal decision-making processes. This inherent lack of clarity poses significant challenges for building trust and verifying the reliability of its outcomes, particularly when applied in sensitive or critical domains like healthcare and autonomous systems. Current literature reveals a multi-faceted approach to interpreting these complex, self-organizing systems. A prominent strategy involves post-hoc explanation methods, where surrogate models—such as decision trees or linear models—are trained on the input-output data generated by the swarm to provide a global, simplified summary of solution features [liang2024survey]. Complementing this, social network analysis (SNA) is used to model the dynamic communication topology of a swarm, quantifying the influence of individual agents and tracing the flow of information that leads to convergence [kennedy2002population]. For a more intrinsic understanding, fitness landscape analysis (FLA) characterizes the problem’s topology to explain algorithmic performance, such as justifying convergence stalls by identifying multimodal ruggedness [malan2013survey]. Furthermore, visual analytics and dimensionality reduction techniques are leveraged to create interactive dashboards and project high-dimensional swarm trajectories into 2D or 3D spaces, allowing users to visually track exploration and exploitation behavior over time [zhang2008sequential]. Despite these advances, challenges remain in standardizing evaluation metrics for explanations and developing real-time, human-in-the-loop XAI frameworks to fully bridge the gap between powerful optimization and user trust. The ambiguous nature of PSO’s configurations and hyperparameters can directly contribute to a perceived lack of reliability in the solutions generated by the algorithm. The stochastic nature of particle movements and the emergent, non-linear interactions within the swarm contribute significantly to this opacity, making it difficult for human users to understand why a specific optimal solution was found or how the swarm navigated the complex search space to converge. This apparent contradiction between PSO’s algorithmic simplicity and its behavioural opacity, reveals a deeper truth: PSO’s algorithmic rules are simple to implement, yet its collective behavior is highly complex and inscrutable. This obscurity does not arise from intricate internal structures but from the swarm’s emergent “intelligence”—the intricate interplay of local and global information sharing across a high-dimensional space. The challenge for XAI is thus that the “black box” nature of such metaheuristics stems from the non-linear dynamics of many simple agents, which confounds a clear tracing of the causal pathway to a solution. Therefore, XAI methods for PSO need to focus on analyzing collective behavior and parameter sensitivity rather than just decomposing complex architectures. In response to this challenge, the fields of Explainable Artificial Intelligence (XAI) and Interpretable Artificial Intelligence (IAI) [doshi2017towards] have rapidly gained prominence. Their primary objective is to develop methodologies that render AI models inherently understandable, thereby clarifying their decision-making processes and nurturing greater confidence in their results.
@@
Historically, metaheuristic optimization algorithms like PSO were primarily evaluated based on their efficiency, convergence speed, and the quality of the solutions they found. However, recent work [gupta2025enhancing] on analysing the impact of swarm topologies provides an early work in the direction of “enhancing explainability and trustworthiness” in PSO. This work is an extension of the work carried out in [gupta2025enhancing]. This signals a significant maturation within the field of computational intelligence. It indicates that achieving optimal solutions is no longer the sole criterion for algorithm success. Instead, there is a growing recognition that for AI systems to be responsibly deployed and widely adopted, especially in critical applications, they must also be transparent and understandable [BARREDOARRIETA202082]. This reflects a broader ethical and practical demand for trustworthy AI. This paradigm shift implies a pressing need for the development of new evaluation metrics that extend beyond traditional optimization performance (e.g., solution quality, computational cost) to quantitatively assess explainability. It also suggests that future research in PSO will increasingly integrate XAI considerations from the very initial design phases of new algorithms, rather than merely applying post-hoc explanations [slack2021reliable]. The follwoing subsections further throws light on some key factors which have the highest impact in the interpretability and explainability of PSO.
@@ 1.1.1 Impact of Communication Topologies in PSO
The effectiveness and search behavior of PSO are profoundly influenced by its communication topology, which defines the structural pattern through which particles share information within the swarm. Common topologies include Ring, [liu2016topology, sun2023particle] Star [miranda2008stochastic, ni2013new], and Von Neumann [von1935complete, lynn2018population]. Research investigates how these distinct topologies impact critical aspects of PSO’s operation, such as information flow, diversity maintenance, convergence speed, and the delicate balance between exploration (searching new regions) and exploitation (refining known good regions) [zhang2019enhancing]. By systematically analyzing these influences, it is possible to enhance the inherent interpretability of PSO’s decision-making processes and derive practical guidelines for selecting the most suitable topology for specific optimization tasks. The explicit detailing of how different communication topologies directly affect PSO’s convergence, diversity, and exploration-exploitation balance provides a powerful form of intrinsic interpretability. This approach is powerful because it moves beyond treating PSO as an indivisible black box. Instead, it focuses on understanding and designing its core components to be inherently transparent. This suggests that modifying the algorithm’s internal architecture and interaction rules can directly contribute to its interpretability, offering a promising avenue for developing more transparent and trustworthy metaheuristics, in future.
@@ 1.1.2 Explainable Benchmarking Frameworks for PSO
The development and application of specialized frameworks, such as the adapted IOHxplainer [van2025explainable], serve as crucial explainable benchmarking tools. These tools can be used to systematically investigate and quantify how different algorithmic configurations, including communication topologies, influence PSO’s performance characteristics. These frameworks are sophisticated enough to support continuous, integer, and categorical parameters, and critically, they account for hyperparameter dependencies, allowing for a detailed analysis of PSO configurations. They often integrate post-hoc XAI methods, such as SHapley Additive exPlanations (SHAP-based techniques) [lundberg2017unified], to determine the marginal impact of each parameter on PSO’s performance, thereby significantly aiding in the interpretability of the experimental results.
@@ 1.1.3 Deeper Understanding Benchmark Problems
Understanding the behaviour of optimization algorithms on benchmark problems is not straightforward, as problem difficulty often depends on hidden structural properties of the search space. Relying only on performance metrics such as convergence speed or solution quality provides limited insight, since these results do not explain why an algorithm succeeds or fails. This is where tools like Exploratory Landscape Analysis (ELA) [mersmann2011exploratory] can play a very crucial role. ELA features allow us to quantify the geometry and structure of optimization problems in a systematic manner. By measuring properties like ruggedness, modality, separability, and variable scaling, ELA helps uncover the reasons behind an algorithm’s performance. For example, highly rugged landscapes may demand more global exploration, while separable problems can be effectively handled by coordinate-wise methods. Such insights are crucial for moving beyond trial-and-error experimentation and towards principled algorithm design and selection. In essence, ELA features have potential to bridge the gap between “black-box” benchmarking and meaningful interpretation. These insights not only guide us in diagnosing problem hardness but also aid in designing adaptive and generalizable optimization strategies.
@@ 1.3 Summary of Contributions
An metaheuristic optimization algorithm’s performance is largely determined by the particular issue it is trying to solve. To figure out why some problems seem tougher for certain algorithms, we look into how algorithms and problems interact. This approach characterizes optimization problems by extracting meta-features from candidate solutions, which are typically generated using sampling techniques such as Latin hypercube sampling  [shields2016generalization], random sampling, or Sobol sequences [burhenne2011sampling]. These sampled solutions form the basis for computing landscape features that capture essential problem properties. For explainable analysis in machine learning, several powerful techniques exist to derive these features, these approaches provide a comprehensive framework for understanding and comparing optimization problems in a meaningful way.
@@
This work extends our earlier work on the explainability of PSO communication topologies [gupta2025enhancing], further enhancing interpretability through the integration of ELA and IOHxplainer-based analysis. The remainder of this paper is organized as follows: Section 2 presents the ELA-based visualization framework employed for benchmarking. Section 3 introduces the proposed explainability framework for PSO, followed by Section 4, which details the experimental setup and corresponding results. The influence of hyperparameters on algorithmic performance is examined in Section 5, while Section 6 outlines the data-driven, interpretable configuration learning methodology. Finally, Section 7 provides the concluding remarks and potential directions for future research.